# Explainable Artificial Intelligence techniques for interpretation of food datasets: a review

### v1

@@ Abstract
Artificial Intelligence (AI) has become essential for analyzing complex data and solving highly-challenging tasks. It is being applied across numerous disciplines beyond computer science, including Food Engineering, where there is a growing demand for accurate and trustworthy predictions to meet stringent food quality standards. However, this requires increasingly complex AI models, raising reliability concerns. In response, eXplainable AI (XAI) has emerged to provide insights into AI decision-making, aiding model interpretation by developers and users. Nevertheless, XAI remains underutilized in Food Engineering, limiting model reliability. For instance, in food quality control, AI models using spectral imaging can detect contaminants or assess freshness levels, but their opaque decision-making process hinders adoption. XAI techniques such as SHAP (Shapley Additive Explanations) and Grad-CAM (Gradient-weighted Class Activation Mapping) can pinpoint which spectral wavelengths or image regions contribute most to a prediction, enhancing transparency and aiding quality control inspectors in verifying AI-generated assessments. This survey presents a taxonomy for classifying food quality research using XAI techniques, organized by data types and explanation methods, to guide researchers in choosing suitable approaches. We also highlight trends, challenges, and opportunities to encourage the adoption of XAI in Food Engineering.
@@ 1 Introduction
Rapid technological advances and the amount of data have made Artificial Intelligence (AI) an essential tool in modern industry and research [1, 2, 3, 4]. Food engineering represents a perfect application for AI technology, as food requires in-depth study, processing, and analysis. The large volume of data generated in this field makes AI especially valuable for data analysis. However, the extensive use of AI introduces new questions about its trustworthiness and reliability.
@@
To ensure trust in the results, it is essential not only to understand the decision-making process behind the AI model but also to enhance its transparency, auditability, and informativeness [5]. Despite this, interpretable AI methods are still not widely adopted in the food sector, highlighting the need for greater focus on transparency and model explainability in this field. In response to this need, eXplainable AI (XAI) has emerged as an important area of research to increase the trustworthiness of AI model predictions. It encompasses techniques aimed at elucidating the behaviour of these models by providing insights into their complex operations. In food engineering, XAI has been applied to allow accurate identification and validation of critical characteristics in tasks such as contaminant detection, nutritional value estimation and product authentication, ensuring safety, transparency and reliability in food quality control. This enables greater confidence by model users and customers, identifies potential biases to improve accuracy, and supports the development of new, safer, and better-quality products.
@@
Given the essential role of food in human life, the food industry is keenly interested in applying these techniques to ensure the reliability of AI-driven outcomes [6]. However, we have identified several gaps in the literature linking XAI with food engineering. Firstly, there is a lack of standardization in the terminology and keywords used across various publications, creating challenges for data analysts and food engineers to communicate effectively. For instance, terms like “interpretation”, “explanation”, and “comprehension” are often used interchangeably for similar tasks, particularly when leveraging AI models in food quality research. Moreover, there is no comprehensive overview of the current state of the art addressing these differences and providing insights about advantages and drawbacks, which could be important to help non-experts understand the progress and potential of these disciplines in research.
@@
This survey provides valuable insights for food industry specialists on the potential and importance of XAI. In particular, it offers an overview of current XAI applications in the food industry across key quality tasks—such as food safety, nutritional value determination, sensory attributes, authenticity and traceability, as well as sustainability and healthiness. We categorize applications by data type (tabular, pictorial, spectral, and time series) and forms of explanations generated by the applied XAI methods (numerical, rule-based, textual, visual, and mixed), highlighting its potential for further development, as depicted in Figure 1. Each task, data type, and form of explainability is discussed in detail in Sections 2, 2.1 and 3.
@@
Additionally, our goal is to bridge the gap between the domains of XAI and food quality by presenting a taxonomy and arranging the current state of XAI applications in food research within an organized structure. Specific objectives are:
@@
Only articles specifically addressing the topic of food quality were considered for this review. Papers that did not explicitly describe the use of a specific “XAI technique” in the field of food quality, even if they covered both food engineering and XAI, were excluded from our analysis. We performed an exhaustive search on Google Scholar and Scopus using the following keywords: “explainable artificial intelligence”, “XAI”, “food”, “food science”, “food quality”, “food control”, and “agriculture”. These terms were strategically combined to cover relevant literature published over the past ten years. We examined the reference sections of the articles obtained in the initial search to identify additional relevant articles and integrated them into our research base. Lastly, we concentrated on several widely used techniques in XAI, including Local Interpretable Model-agnostic Explanations (LIME) [7], SHapley Additive exPlanations (SHAP) [8], Class Activation Mapping (CAM) [9], Partial Dependence Plot (PDP) [10], and Layer-wise Relevance Propagation (LRP) [11]. We investigated papers citing these foundational works to identify any additional relevant articles and incorporated them into our research base.
@@ 2.1 Data Types
With the continuous advancement of technology, food quality analysis has significantly evolved, leveraging the diversity of sensors, methods and devices to collect data into datasets. These datasets encompass various modalities, including tabular data, images or pictorial data, spectral data, and time series data, each offering distinct advantages for analysts in evaluating crucial aspects of food quality. The complexity and volume of these data have necessitated AI to process large datasets automatically and identify complex patterns, extracting the maximum useful information from these diverse data.
@@ 3 XAI methods
Explainability and interpretability in the context of AI models, despite often being used interchangeably, the two notions are slightly different in meaning, as explained by [42]. The authors argue that interpretability is concerned with understanding the inner workings of a model, while explainability is strictly tied to providing post-hoc, approximate insights on a prediction operated by the model. In other words, interpretability is an intrinsic property of a model, while an explanation is generated on a (non-)interpretable model after a prediction has been made.
@@
Accuracy vs. interpretability trade-off: Interpretability, as defined above, has often been depicted as being at odds with expressivity or accuracy [^1] In this specific case, we use the term accuracy as a generic stand-in for the performance of the model in solving the task which it was designed to carry out. of the model [43]. Expressivity, also termed flexibility, refers to the range of complicated patterns that the model can learn. LR is often depicted as a very inflexible model since it can only learn simple linear relationships between predictors; hence, its accuracy will be fairly limited on more complex problems, like those linked to pictorial data. However, the linear relationship is interpretable by human standards: a single parameter of an LR model indicates the additive effect that a perturbation of the corresponding predictor has on the response. This makes it straightforward to analyze, for instance, the importance of each variable within the model.
@@
On the other hand, highly expressive models like Deep NNs are seen as complex and are hence usually tagged as black-boxes. Indeed, despite reaching high accuracy on very complex problems, it is often hard to gain an interpretation of the rules learned by such models for achieving a certain prediction. A depiction of this trade-off can be seen in Figure 2, where AI models from the analyzed studies, as described in the Section 2.2, are positioned accordingly. Despite the trade-off being renowned in the literature, it is still an approximate rule-of-thumb, which has exceptions, like in the case of vision transformers [44], which, despite being more expressive than CNNs, are defined as inherently more interpretable due to the ease of visualizing the attention mechanism [45].
@@
Figure 2: Chart illustrating approximately the trade-off between expressivity or flexibility and interpretability. Expressive models, such as those based on DL, are capable of reaching higher task-level performance but are often hardly interpretable. On the other hand, less complex models, like LR, are inherently interpretable, but often incapable of attaining high task-level performance.
@@
Global vs. local XAI methods: Another axis which defines XAI tools is represented by the scope of the method. If the tool delves into properties of the model as a whole, then the scope is said to be global; conversely, when the tool investigates the model behavior around one data point, then the scope is said to be local. Concerning the LR example before, the model’s coefficients can be thought of as global explanations, since they define a global behaviour of the model irrespective of the specific data point considered. On the other hand, as an example of local explanation, we can consider feature attribution in the context of image classification using CNNs. For feature attribution, we indicate the action of identifying which variables contribute the most to producing the prediction. In the case of image classification, it may be of interest to elicit important pixels that led a given picture to be classified in a given category; this is an example of a local explanation since we are gaining knowledge of the behaviuor of the model only on the current image, without trying to infer the global properties of the model. In the case of NNs, it is often hard to identify such global rules for explaining predictions; thus, local explanations are often preferred [46]. Despite being limited in scope, local explanations can be used to extrapolate global information about the models, as, for instance, in the works by [46] and [47].
@@
Model-agnostic vs. model-specific XAI methods: A final property of the XAI tools to be considered is the specificity to limited classes of models. Model-agnostic tools are XAI methods that, due to how they are constructed, can be applied to any AI model, while model-specific tools are restricted to limited classes of models. In the aforementioned case of feature attribution, methods like LIME and SHAP are considered model-agnostic; on the other hand, techniques like CAM and Grad-CAM [48] are applicable only in case of CNNs for classification, although the latter has been extended to other NN architectures, like Vision Transformers [49], or to other tasks, like regression [50].

### v2

# Explainable Artificial Intelligence Techniques for Interpretation of Food Models: a Review

@@ Abstract
Artificial Intelligence (AI) has become essential for analyzing complex data and solving highly-challenging tasks. It is being applied across numerous disciplines beyond computer science, including Food Engineering, where there is a growing demand for accurate and reliable predictions to meet stringent food quality standards. However, this requires increasingly complex AI models, raising concerns. In response, eXplainable AI (XAI) has emerged to provide insights into AI decision-making, aiding model interpretation by developers and users. Nevertheless, XAI remains underutilized in Food Engineering, limiting model reliability. For instance, in food quality control, AI models using spectral imaging can detect contaminants or assess freshness levels, but their opaque decision-making process hinders adoption. XAI techniques such as SHAP (Shapley Additive Explanations) and Grad-CAM (Gradient-weighted Class Activation Mapping) can pinpoint which spectral wavelengths or image regions contribute most to a prediction, enhancing transparency and aiding quality control inspectors in verifying AI-generated assessments. This survey presents a taxonomy for classifying food quality research using XAI techniques, organized by data types and explanation methods, to guide researchers in choosing suitable approaches. We also highlight trends, challenges, and opportunities to encourage the adoption of XAI in Food Engineering.
@@ 1 Introduction
Rapid technological advances and the amount of data have made Artificial Intelligence (AI) an essential tool in modern industry and research [100, 23, 232, 171]. Food engineering represents a perfect application for AI technology, as food requires in-depth study, processing, and analysis. The large volume of data generated in this field makes AI especially valuable for data analysis. However, the extensive use of AI introduces new questions about its reliability.
@@
To ensure reliable results, it is essential not only to understand the decision-making process behind the AI model but also to enhance its transparency, auditability, and informativeness [124]. Despite this, interpretable AI methods are still not widely adopted in the food sector, highlighting the need for greater focus on model transparency in this field. In response to this need, eXplainable AI (XAI) has emerged as an important area of research to increase the reliability of AI model predictions. It encompasses techniques aimed at elucidating the behaviour of these models by providing insights into their complex operations. In food engineering, XAI applications allow accurate identification and validation of critical characteristics in tasks such as contaminant detection, nutritional value estimation, and product authentication, ensuring safety, transparency, and reliability in food quality control. This enables greater confidence by model users and customers, identifies potential biases to improve accuracy, and supports the development of new, safer, and better-quality products.
@@
Given the essential role of food in human life, the food industry is keenly interested in applying these techniques to ensure the reliability of AI-driven outcomes [137]. However, we have identified several gaps in the literature linking XAI with food engineering. Firstly, there is a lack of standardization in the terminology and keywords used across various publications, creating challenges for data analysts and food engineers to communicate effectively. For instance, terms like “interpretation”, “explanation”, and “comprehension” are often used interchangeably for similar tasks, particularly when leveraging AI models in food quality research. Moreover, there is no comprehensive overview of the current state of the art addressing these differences and providing insights about advantages and drawbacks, which could be important to help non-experts understand the progress and potential of these disciplines in research. Existing literature already includes several surveys collecting works that compile and organize research on XAI. However, these contributions address different themes and do not specifically focus on the application of XAI to the food industry. Some present XAI in a purely theoretical way, without considering concrete use cases [21, 124, 149]. Additionally, there are surveys on AI applications that do not directly address explainability [210, 204], as well as narrow studies, such as those focused on fruit [278] or plant leaves [246], that are limited to specific contexts.
To address these gaps and respond to the growing need to build reliability in AI outcomes through modern XAI, we propose this survey, which offers a comprehensive overview of XAI applications in food engineering, with explainability explicitly placed at the core of the analysis. This survey is primarily intended for researchers, practitioners, and professionals in food science, food industry specialists, and food engineers, who can gain a clear view of the state of the art in XAI applications within the food sector and understand their relevance. Additionally, this survey serves as a reference guide for AI developers and policymakers who are involved in designing or evaluating explainable systems related to food safety, quality control, and traceability. By presenting various techniques and their applications in different contexts, the survey helps readers easily identify available methods and facilitates the acquisition of new knowledge in the field.
In particular, it offers an overview of current XAI applications in the food industry across major quality-related tasks, such as food safety, nutritional value determination, sensory attributes, authenticity and traceability, sustainability, and healthiness.
@@
To provide a clear reading framework, we categorize applications by data type (tabular, pictorial, spectral, and time series) and forms of explanations generated by the applied XAI methods (numerical, rule-based, textual, visual, and mixed), highlighting their potential for further development, as depicted in Figure 1. Each task, data type, and form of explainability is discussed in detail in Sections 2, 2.2, and 3, respectively.
@@
Additionally, we aim to bridge the gap between XAI and food quality research by proposing a taxonomy and systematically organizing the current landscape of XAI applications in food-related studies. Key contributions of this paper are:
@@
1.1 Review methodology
We conducted a semi-systematic literature review, combining structured database queries with backward and forward snowballing to improve coverage in a fast-evolving interdisciplinary area. We queried Scopus and Google Scholar using combinations of the keywords: “explainable artificial intelligence”, “XAI”, “food”, “food science”, “food quality”, “food control”, and “agriculture”. We focused on works published in the last ten years and included studies available up to July 2025. We applied a targeted eligibility criterion: we retained only studies explicitly addressing food quality and explicitly employing at least one XAI technique; papers that discussed AI and food but without an explicit XAI method were excluded. Finally, we expanded the initial corpus through backward screening of reference lists and forward citation chasing, starting from foundational XAI techniques (e.g., LIME, SHAP, CAM/Grad-CAM, PDP, LRP). Figure 2 summarizes the workflow.
Identification
Databases:
Scopus; Google Scholar
Search terms: “explainable artificial intelligence”, “XAI”, “food”, “food science”, “food quality”, “food control”, “agriculture” (combined) Time window: last 10 years; works available up to July 2025
Screening
Title/abstract screening for relevance to food quality and XAI
Eligibility
Full-text assessment
Include: explicit food quality focus + explicit XAI technique Exclude: food-related AI without explicit XAI method
Snowballing (iterative)
Backward: references of eligible papers
Forward: citation chasing from foundational XAI techniques (LIME, SHAP, CAM/Grad-CAM, PDP, LRP)
Included studies
Final corpus: >>200 works surveyed
Figure 2: Semi-systematic literature review workflow. Structured queries on Scopus and Google Scholar using XAI and food-related keywords were combined with backward and forward snowballing from foundational XAI techniques (e.g., LIME, SHAP, Grad-CAM). Studies were retained only if they explicitly addressed food quality and employed at least one XAI method.
1.1.1 Methodological limitations
This survey is semi-systematic rather than fully systematic: while structured searches were performed, the final corpus also depends on iterative snowballing, which may introduce selection bias toward highly cited works. Second, restricting the search to Scopus and Google Scholar can omit relevant studies indexed in other databases. Third, keyword-based retrieval may miss papers that adopt non-standard terminology for explainability or food quality tasks. Finally, our coverage is time-bounded (up to July 2025) and thus does not include subsequent publications.
@@ 2 Explaining Food Quality
2.1 Food-industry context and domain-specific challenges
The tasks examined in the previous section represent the primary operations routinely addressed by food quality analysis within the food supply chain. However, these tasks present domain-specific challenges that substantially influence both model development and the interpretation of model behavior. Unlike many controlled laboratory settings, industrial food pipelines are characterized by biological variability arising from genetics, physiology, and microbiota, as well as heterogeneity in raw materials with respect to size, composition, water activity, and fat-to-protein ratio. Additional sources of variation include seasonal effects driven by feed composition, climate, and cultivation conditions, alongside batch-to-batch fluctuations attributable to sourcing and production parameters. Collectively, these factors frequently induce distribution shifts across time, production sites, and suppliers, thereby increasing the risk that models exploit spurious correlations that fail to generalize beyond the training conditions.
In addition, food processing and storage introduce confounding factors. Changes in temperature profiles, packaging type and permeability, oxygen exposure, humidity, and storage duration can alter sensory and physicochemical properties in ways that correlate with the target outcome but are not causally linked to the underlying quality attribute. For example, a vision model trained for defect detection may inadvertently learn to rely on acquisition artifacts such as lighting conditions, background appearance, or conveyor patterns. Similarly, a spectral model developed for freshness estimation may capture instrument drift, preprocessing choices, or incidentally correlated wavelength regions rather than robust chemical signatures. These realities motivate the adoption of XAI as a diagnostic layer to support verification, auditing, and root-cause analysis of model predictions under operational conditions.
Illustrative industry scenarios.
To illustrate the practical relevance of XAI, we present a series of concise scenarios that frequently arise in food industry practice, each highlighting contexts in which XAI can meaningfully support decision-making in realistic operational settings.
(S1) Contamination detection through imaging-based inspection. In high-throughput inspection settings, such as foreign-body detection, surface contamination assessment, and visible defect identification, XAI can assist inspectors in verifying whether a model focuses on plausible image regions, such as the contaminant area itself, rather than background cues, including packaging elements or lighting reflections [130, 237, 224]. This capability supports rapid triage, facilitates communication between operators and data scientists, and enables targeted re-acquisition or sensor recalibration when explanations systematically highlight non-causal artifacts as the primary drivers of model predictions [151, 160, 104].
(S2) Freshness assessment and shelf-life monitoring across multimodal and spectral data. In freshness prediction tasks that rely on spectral measurements or on combined sensor streams integrating temperature and humidity readings with product-level measurements, feature attribution methods can identify which spectral bands or environmental variables are responsible for triggering a quality alarm [213, 227, 37]. Quality managers can leverage these explanations to assess whether a given prediction aligns with known degradation mechanisms, and to determine whether corrective actions should target process parameters such as cooling-chain control or instead focus on product handling procedures.
(S3) Authenticity verification and traceability through tabular and spectral fingerprints. In authenticity verification tasks such as geographical origin discrimination and adulteration detection, attribution-based methods can highlight the most influential chemical markers or elemental fingerprints driving a classification decision [18, 220]. These explanations support auditing processes by documenting which indicators contributed to a given outcome, thereby helping laboratories and regulatory compliance teams justify their conclusions and prioritize the selection of confirmatory analyses [2, 172, 97, 85].
(S4) Quality grading and process optimization from tabular process variables. In grading tasks governed by process and physicochemical variables, explainability tools can reveal population-level trends, such as the influence of temperature or processing time on predicted grade, while also helping operators interpret specific borderline cases. This fosters an actionable understanding of model behavior and supports the definition of standard operating procedures for targeted interventions, as illustrated by the tabular food quality applications surveyed in this work, including quality grading and physicochemical-driven assessments in beverage and dairy production contexts [118, 201, 176, 261, 260].
@@ 2.2 Data Types
With the continuous advancement of technology, food quality analysis has significantly evolved, leveraging the diversity of sensors, methods, and devices to collect data into datasets. These datasets encompass various modalities, including tabular data, images or pictorial data, spectral data, and time series data, each offering distinct advantages for analysts in evaluating crucial aspects of food quality. The complexity and volume of these data have necessitated AI to process large datasets automatically and identify complex patterns, extracting the maximum useful information from these diverse data.
@@ 3 XAI methods
As previously mentioned, XAI revolves around the task of explaining (parts of) a model to human stakeholders in an effort to improve the reliability or the transparency of an AI system. Its outputs are termed explanations. This task is usually split in two sub-categories, explainability and interpretability, depending on the approach adopted.
@@
Explainability vs. Interpretability: despite often being used interchangeably, these two notions are slightly different in meaning, as explained by [31]. The author argues that interpretability is concerned with understanding the inner workings of a model, while explainability is strictly tied to providing post-hoc, approximate insights on a prediction operated by the model. In other words, interpretability can be seen as an intrinsic property of a model, while an explanation has a post-hoc meaning: it is strictly generated on a (non-)interpretable model after the model has been trained.
We provide a glossary table with this distinction in Table 1. We additionally point out that, in the context of the present work, we will mostly refrain from using terms such as “understandability” or “comprehensibility”, since they do not appear to be part of a formal XAI taxonomy, and “trust” and “transparency”, since their meaning extends far beyond the scope of XAI.
Table 1: Key taxonomy in XAI: interpretability as an intrinsic model property and explainability as a post-hoc analysis tool.
Interpretability Intrinsic property of the model: degree over which the predictive logics of the model are naturally understandable to a human.
Main interpretable models: LR, decision trees, fuzzy logic.
Explainability Creation of explanations after (post-hoc) a model has been trained: they provide (approximate) insights over (a part of) a (non-interpretable) model to make the prediction itself more interpretable to a human. An overview of the main explainability methods is given in Section 3.2.
@@
Accuracy vs. interpretability trade-off: Interpretability, as defined above, has often been depicted as being at odds with accuracy [^1] In this specific case, we use the term accuracy as a generic stand-in for the performance of the model in solving the task which it was designed to carry out., also termed expressivity or flexibility of the model [94]. Flexibility refers to the range of complicated patterns that the model can learn. LR is often depicted as a very inflexible model because it can only learn simple linear relationships between predictors; hence, its accuracy will be fairly limited on more complex problems, such as those involving pictorial data. However, the linear relationship is interpretable by human standards: a single parameter of an LR model indicates the additive effect that a perturbation of the corresponding predictor has on the response. This makes it straightforward to analyze, for instance, the importance of each variable within the model.
@@
Conversely, highly expressive models such as Deep NNs are considered complex. While they achieve high accuracy on very intricate problems, interpreting the latent rules these models learn to make specific predictions can be quite challenging. A depiction of this trade-off can be seen in Figure 3, where AI models from the analyzed studies, as described in the Section 2.3, are positioned accordingly. Despite the trade-off being renowned in the literature, it is still an approximate rule-of-thumb, which has exceptions, like in the case of vision transformers [59], which, despite being more expressive than CNNs, are defined as inherently more interpretable due to the ease of visualizing the attention mechanism [211].
@@
Figure 3: The trade-off between expressivity or flexibility and interpretability of the AI models exploited in the reviewed studies. Expressive models, such as those based on DL, are capable of reaching higher task-level performance but are often difficult to interpret. On the other hand, less complex models, like LR, are inherently interpretable, but often incapable of attaining high task-level performance.
@@
Global vs. local XAI methods: Another axis that defines XAI tools is represented by the scope of the method. If the tool delves into properties of the model as a whole, then the scope is said to be global; conversely, when the tool investigates the model behavior around one data point, then the scope is said to be local. Concerning the LR example before, the model’s coefficients can be thought of as global explanations, since they define a global behaviour of the model irrespective of the specific data point considered. On the other hand, as an example of local explanation, we can consider feature attribution in the context of image classification using CNNs. For feature attribution, we indicate the action of identifying which variables contribute the most to producing the prediction. In the case of image classification, it may be of interest to elicit important pixels that led a given picture to be classified in a given category; this is an example of a local explanation since we are gaining knowledge of the behavior of the model only on the current image, without trying to infer global properties. In the case of NNs, it is often hard to identify such global rules for explaining predictions; thus, local explanations are often preferred [253]. Despite being limited in scope, local explanations can be used to extrapolate global information about the models [253, 202].
@@
Model-agnostic vs. model-specific XAI methods: A final property of the XAI tools to be considered is the specificity to limited classes of models. Model-agnostic tools are XAI methods that, due to how they are constructed, can be applied to any AI model, while model-specific tools are restricted to limited classes of models.
@@ 3.1 Explanation Types
3.2 Overview of Popular Explainable Artificial Intelligence Techniques
In recent years, several explainability techniques have gained importance in response to the growing need for transparent AI models. These methods are generally straightforward to apply and support a range of models and explanation types. Together, they constitute a versatile toolkit that can be adapted to diverse cases and operational contexts. In this Section, we do not detail every technique because of space constraints. Instead, we just concentrate on the main model-agnostic and model-specific methods, as defined by [241]. We present a concise summary that highlights the key aspects, indicates the explanation type for each method, and provides a general classification based on the taxonomy introduced in the preceding section. An overview of these methods is also reported in Table 2.
LIME [191] explains a single prediction by training an interpretable surrogate model. This model is trained to mimic the black box model in the neighborhood of the instance to explain. It involves perturbing the instance, creating $k$ synthetic data points, querying the model on these nearby samples, and weighting them based on their proximity. The interpretable model is usually a LR or a small decision tree. The coefficients of the LR, or the feature importance estimates of the decision tree are used as proxies for the feature importance of the black box model. These values are numerical explanations, but they can also be plotted (e.g., as a bar plot or a heatmap) to create a visual representation.
SHAP [127] explains a single prediction by attributing to each feature its game-theoretic contribution to the model’s output. It compares the prediction for the instance with predictions with one or more features removed, then aggregates each feature’s marginal effect across feature combinations (coalitions) to compute its Shapley value. The feature removal is simulated by replacing the values of the features with a reference baseline (e.g., the average value of each feature). The explanation is additive, so the sum of feature contributions equals the difference between the model’s expected output and the instance’s prediction. While the explanation is numerical, it is typically visualized with dedicated attribution plots. It can also be visualized as a heatmap, e.g., to explain visual data.
PDP [70] explains model behavior by showing the marginal effect of one or two features on the predicted outcome. They are computed by varying the selected feature values over a grid, querying the model at each value while averaging predictions over the distribution of all other features, and assembling the resulting partial dependence function. This isolates the average relationship between the chosen features and the prediction, revealing trends such as monotonicity, saturation, or interaction strength. The explanation consists of a line plot for one feature or a surface plot for two features.
LRP [16], CAM [275], and its variant Grad-CAM [200] are NN–specific tools. They explain a single prediction by tracing class evidence back to the input. LRP backpropagates the output score through the network with relevance conservation, redistributing relevance at each layer in proportion to local contributions to yield a feature relevance map. CAM and Grad-CAM compute a class-specific localization map by weighting the final feature maps with the gradients of the class weights and aggregating them to highlight class-discriminative regions. CAM and Grad-CAM are originally designed for CNNs for classification trained on image data, although there exist extensions for other architectures, tasks, and data types. Both methods generate numerical attributions that are typically represented as heatmaps overlaid on the input.
Table 2: Overview of the most popular XAI methods classified by scope, specificity, and explanation type. Efficiency is an indication of the computational complexity ($\mathcal{O}(\cdot)$) for running the method on a single data point, on which we want to generate an explanation. $d$ indicates the number of features, while $k$ indicates the number of additional data points created via perturbation (for LIME).
XAI technique Scope Specificity Explanation type Efficiency
LIME [191] Local Model-agnostic Numerical / Mixed $\mathcal{O}(kd^{2})$
SHAP [127] Local Model-agnostic Numerical / Mixed $\mathcal{O}(d2^{d})$
PDP [70] Global Model-agnostic Visual $\mathcal{O}(d)$
LRP [16] Local Model-specific Visual $\mathcal{O}(1)$
CAM [275] Local Model-specific Visual $\mathcal{O}(1)$
Grad-CAM [200] Local Model-specific Visual $\mathcal{O}(1)$
Figure 5: Examples of the main techniques introduced in Section 3.2. LIME (numerical output shown in vertical bar plots), SHAP (numerical output shown as a force plot, and heatmap for image data), PDP (line plots), LRP (heatmap), and CAM/Grad-CAM (heatmap). Diagrams are adapted from the following references: LIME, [98]; SHAP, [233] and [221]; PDP, [65]; LRP, [226]; CAM/Grad-CAM, [126].
Overall, Table 2 highlights how popular model-specific techniques (LRP, CAM, Grad-CAM) have a constant computational complexity with respect to the input dimension; on the other hand, model-agnostic tools, despite their flexibility, all have at least a linear complexity. This provides model-specific methods with a clear efficiency advantage and motivates the widespread adoption of such methods observed in several of the papers surveyed in the present work. Additionally, in Figure 5, we provide an example of visualization for each of the techniques detailed in this Section. When comparing model-specific tools, most of the works hereby surveyed prefer the usage of CAM-based methods instead of LRP. It has indeed been observed empirically that human participants tend to find the coarser-grained explanations produced by Grad-CAM more interpretable [109, 104]. This justifies the wider adoption of Grad-CAM with regards to LRP, despite similar efficiencies.